OpenStudy (anonymous):

Find the exact value of the radical expression in simplest form. the square root of 16 times y cubed − the square root of y cubed?

OpenStudy (anonymous):

please help me! This is veryyyy important:))))

OpenStudy (dumbcow):

\[\sqrt{16y^{3}} - \sqrt{y^{3}} = \sqrt{16}\sqrt{y^{3}} - \sqrt{y^{3}} = 4\sqrt{y^{3}}-\sqrt{y^{3}} = 3\sqrt{y^{3}}\]

OpenStudy (anonymous):

thank you dumbcow:)

OpenStudy (dumbcow):

Also \[\sqrt{y^{3}} = \sqrt{y^{2}}\sqrt{y} = y \sqrt{y}\]